# The Atacama Cosmology Telescope: The Development of Machine Learning Tools for Detecting Millimeter Sources in Timestream Pre-processing

### v1

@@ 8 Conclusions
Our improved data cuts algorithm enables classifications of glitches into point sources, point sources with another coincident glitch, cosmic rays, and electronic glitches from the output of the ACT glitch finder. We have an overall accuracy of 90%, with 94% of PS objects being correctly classified. We are able to restore high-amplitude sources that previously would have been removed from the maps and flag the sources that would have partially passed the cuts. This enables us not only to ensure that sources are not removed from the maps but also to generate a list of times and positions of transients. Such a list can be used to issue rapid alerts and also provide informative priors for map-based searches.
@@
The algorithm selects sources that appear multiple times in a TOD. The amplitudes of such sources can be computed across scans to generate sub-minute to minute timescale light curves. We tested the recovery of simulated light curves of the stellar flares using this pipeline. TOD-specific systematics were averaged out in order to better understand the performance we could generally expect from the amplitude computation. We found that for a bright 12,000 mJy source we can detect variability with for a flare with half-life s, and for a dimmer, 6000 mJy source the significance was 1.5–2 for half-lives between 100 s and 500 s, with no detection of longer or shorter flares. These results are from ACT PA4 and PA5, and with more sensitive instruments, like the upcoming SO 3, 2, the ability to determine whether a source is varying on short time scales will improve. Finally, even when a candidate transient has low SNR, or is only seen in one or two scans, we can perform forced-aperture photometry on the source location directly in the TOD to obtain a longer light curve with a baseline flux before the transient appears.
@@
We are adapting our algorithm for SO. In addition to providing data cuts for the main map-making pipeline, its transient detection capabilities are highly complementary to the map-based transient searches which are being developed. In addition to the benefits mentioned above—preventing transients from being excised from the data and providing a list of transient candidates to map-based searches—it can create high time-resolution light curves (sub-minute to minute) of transients, which cannot be readily extracted from maps. This will be helpful in issuing faster transient alerts to the broader community, since current ACT map-based transient searches require multiple TODs and the creation of depth-1 maps. Additionally, the light curves retrieved using this algorithm could allow us to rapidly classify the potential transient type (e.g. stellar flares, which are short duration, versus extragalactic explosions like gamma ray bursts, which are long duration). Furthermore, our experience thus far indicates that TOD searches may perform better on bright transients in the Galactic plane where the source density is higher, and it is more challenging for map-based searches to disentangle nearby sources from each other.
@@
Other avenues for future work include, for instance, adding more statistics to reduce misclassified glitches, particularly for a point source with another spatially coincident glitch in the focal plane. Another interesting avenue would be to create a finer electronic glitch categorization to aid in the understanding of telescope systematics. This would especially help diagnose issues that arise when building a new telescope or implementing new hardware. This could be achieved by first implementing an unsupervised clustering algorithm to determine which categories of electronic glitches arise. We focused on using summary statistics in order to have a light-weight algorithm that can be run in real time during pre-processing. Other approaches, even though they are more memory intensive, such as image classification on the focal plane and timestreams, should also be explored. Another important area for future development is to improve the initial glitch detection algorithm so that it has a lower cut-off, since the relatively high threshold currently being used is a limiting factor in our performance. We are currently developing methods using machine learning to better detect glitches and sources in the data, with the goal of integrating detection into the classification algorithm. Finally, we have demonstrated our ability to recover sub-minute light curves directly from the TODs, but our simple fitting method could be further improved, particularly to obtain better error estimates for the light curves.
@@ Appendix A Classification Results with an Ideal Cuts Algorithm
The current cuts algorithm requires an cutoff of 10 which results in only bright sources to be flagged. In order to test if our algorithm would be able to classify dimmer sources if they were flagged by the cuts, we utilized the stellar flare simulations to create “ideal cuts”. Since the sources are simulated, we know the amplitudes at all times. We create modified cuts objects that flag every detector at a given time that has an amplitude greater than 500 mJy, chosen to be just below the average noise of the detectors. We ran these modified cuts through our classification algorithm, and the results can be seen in Figure 18. For all half-lives and amplitudes, we detected the source using these ideal cuts. This is very promising and is likely due to the importance of the focal plane statistics to our classification algorithm, as we see Figure 7. In general, we see a decrease in probability for later scans as the source’s amplitude is decaying following Equation 6. However, this trend is not strictly followed for the first couple scans, as the ACT focal plane is a hexagon oriented such that the scan begins at one of its vertices. Thus, when the source first enters the focal plane, it only hits a few detectors. As the sky rotates through the focal plane, which is scanning at constant elevation, the source moves towards its center and hits more detectors; finally, as it exits the focal plane it again only hits a few detectors. We note that for 3000 mJy, fewer than four detectors were flagged for the first three scans and thus were not classified, as we require 4 detectors for classification.

### v2

@@ 8 Conclusions
Our improved data cuts classification pipeline enables classifications of glitches into point sources, point sources with another coincident glitch, cosmic rays, and electronic glitches from the output of the ACT glitch finder. We have an overall accuracy of 90%, with 94% of PS objects being correctly classified. We are able to restore high-amplitude sources that previously would have been removed from the maps and flag the sources that would have partially passed the cuts. This enables us not only to ensure that sources are not removed from the maps but also to generate a list of times and positions of transients. Such a list can be used to issue rapid alerts and also provide informative priors for map-based searches.
@@
The pipeline selects sources that appear multiple times in a TOD. The amplitudes of such sources can be computed across scans to generate sub-minute to minute timescale light curves. We tested the recovery of simulated light curves of the stellar flares using this pipeline. TOD-specific systematics were averaged out in order to better understand the performance we could generally expect from the amplitude computation. We found that for a bright 12,000 mJy source we can detect variability with for a flare with half-life s, and for a dimmer, 6000 mJy source the significance was 1.5–2 for half-lives between 100 s and 500 s, with no detection of longer or shorter flares. These results are from ACT PA4 and PA5, and with more sensitive instruments, like the upcoming SO 3, 2, the ability to determine whether a source is varying on short time scales will improve. Finally, even when a candidate transient has low SNR, or is only seen in one or two scans, we can perform forced-aperture photometry on the source location directly in the TOD to obtain a longer light curve with a baseline flux before the transient appears.
@@
We are adapting our classification pipeline for SO. In addition to providing data cuts for the main map-making pipeline, its transient detection capabilities are highly complementary to the map-based transient searches which are being developed. In addition to the benefits mentioned above—preventing transients from being excised from the data and providing a list of transient candidates to map-based searches—it can create high time-resolution light curves (sub-minute to minute) of transients, which cannot be readily extracted from maps. This will be helpful in issuing faster transient alerts to the broader community, since current ACT map-based transient searches require multiple TODs and the creation of depth-1 maps. Additionally, the light curves retrieved using this pipeline could allow us to rapidly classify the potential transient type (e.g. stellar flares, which are short duration, versus extragalactic explosions like GRBs, which are long duration). Furthermore, our experience thus far indicates that TOD searches may perform better on bright transients in the Galactic plane where the source density is higher, and it is more challenging for map-based searches to disentangle nearby sources from each other.
@@
Other avenues for future work include, for instance, adding more statistics to reduce misclassified glitches, particularly for a point source with another spatially coincident glitch in the focal plane. Another interesting avenue would be to create a finer electronic glitch categorization to aid in the understanding of telescope systematics. This would especially help diagnose issues that arise when building a new telescope or implementing new hardware. This could be achieved by first implementing an unsupervised clustering algorithm to determine which categories of electronic glitches arise. We focused on using summary statistics in order to have a light-weight pipeline that can be run in real time during pre-processing. Other approaches, even though they are more memory intensive, such as image classification on the focal plane and timestreams, should also be explored. Another important area for future development is to improve the initial glitch detection algorithm so that it has a lower cut-off, since the relatively high threshold currently being used is a limiting factor in our performance. We are currently developing methods using machine learning to better detect glitches and sources in the data, with the goal of integrating detection into the classification pipeline. Finally, we have demonstrated our ability to recover sub-minute light curves directly from the TODs, but our simple fitting method could be further improved, particularly to obtain better error estimates for the light curves.
@@ Appendix A Classification Results with an Ideal Cuts Algorithm
The current cuts algorithm requires an cutoff of 10 which results in only bright sources to be flagged. In order to test if our trained random forest would be able to classify dimmer sources if they were flagged by the cuts, we utilized the stellar flare simulations to create “ideal cuts”. Since the sources are simulated, we know the amplitudes at all times. We create modified cuts objects that flag every detector at a given time that has an amplitude greater than 500 mJy, chosen to be just below the average noise of the detectors. We ran these modified cuts through our classification pipeline, and the results can be seen in Figure 18. For all half-lives and amplitudes, we detected the source using these ideal cuts. This is very promising and is likely due to the importance of the focal plane statistics to our classification pipeline, as we see Figure 10. In general, we see a decrease in probability for later scans as the source’s amplitude is decaying following Equation 6. However, this trend is not strictly followed for the first couple scans, as the ACT focal plane is a hexagon oriented such that the scan begins at one of its vertices. Thus, when the source first enters the focal plane, it only hits a few detectors. As the sky rotates through the focal plane, which is scanning at constant elevation, the source moves towards its center and hits more detectors; finally, as it exits the focal plane it again only hits a few detectors. We note that for 3000 mJy, fewer than four detectors were flagged for the first three scans and thus were not classified, as we require 4 detectors for classification.